a beach ball is inflated to a volume of 25L of air at 15 degree Celsius. during the afternoon, the volume increases by 1L. what is the new temperature outside?


1
Expert's answer
2021-10-13T04:47:06-0400

Since the amount of gass is fixed, Charles law will apply. Volume is directly proportional to absolute temperature.

"\frac{Volume 1}{Temperature 1}=\frac{Volume 2}{Temperature 2}= constant"

Volume 1=25L Volume 2=25+1=26L

Temperature 1=273+15=288K

Temperature 2 will be "\frac{volume 2 × temperature 1}{volume 1}=\frac{26×288}{25}=299.52K\space\space Or\space299.52-273=26.52°C"